need help solving a problem
-4x-6y+15=0

Answer by JBarnum(2146)

-4x-6y+15=0
-4x=6y-15
x=(-6y+15)/4 for example if y=-1 (6+15)/4= 21/4= 5.25=x
-6y=4x-15
y=(-4x+15)/6